8. Air trapped in a cylinder fitted with a piston occupies 145.7 mL at 1.08 atm pressure. What is the new volume of air when the pressure is increased to 1.43 atm by applying force to the piston? (110. mL)


1
Expert's answer
2021-04-21T03:54:01-0400

P1V1=P2V2

V2 = P1V1/P2 = 145.7 × 1.08 / 1.43 = 110mL